OpenStudy (anonymous):

If I have SinA and CosB how do I find Cos(A+B) and Sin(A+B)? (Preferably without a calculator)

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

use the "addition angle formula" but you need to find two other numbers first

OpenStudy (misty1212):

you need \(\cos(A)\) and \(\sin(B)\) do you know how to find them?

OpenStudy (anonymous):

Example: Sin(a)=5/13 Cos(b)=4/5 Find the value of sin(a+b)

OpenStudy (misty1212):

if \(\sin(A)=\frac{5}{13}\) then \(\cos(A)=\frac{12}{13}\) assuming A is in quadrant 1

OpenStudy (anonymous):

Yes I apologize I forgot to mention that they are acute angles in the first quadrant